Circular No. 5878

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

X-RAY NOVAE A. Kaniovsky, K. Borozdin, and R. Sunyaev, on behalf of the Mir-

Kvant team, report: “The soft x-ray transient GRS 1009–45, discovered by Granat/WATCH and more precisely localized by BATSE (IAUC 5864), was observed by the TTM and HEXE instruments of Mir-Kvant on Sept.

30. Preliminary analysis shows that the source spectrum in the band 2-200 keV is similar to the spectra of the x-ray novae QZ Vul = GS 2000+25 and GU Mus (= Nova Mus 1991) = GRS 1124–684. A bright, soft component can be approximated by a black-body spectrum with a temperature of 0.52 ± 0.03 keV. The source spectrum at the higher energies 10-100 keV is well approximated by a power law with spectral index –2.53 ± 0.05. The flux from the source at 3 and 20 keV was equal to 1 and 0.1 Crab unit, respectively.”

K. Borozdin, V. Arefiev, and R. Sunyaev, on behalf of the Mir-Kvant-TTM team, report: “Mir-Kvant observed the hard x-ray transient GRS 1716–249 = GRO J1719–24 and found that the source spectrum in the band 2-27 keV is approximated by a single power law with spectral index −1.57 ± 0.13. No trails of the soft component were detected in the spectrum; it is reminiscent of the spectra of V1357 Cyg = Cygnus X-1, 1E 1740.7 − 2942, and the x-ray novae V404 Cyg = GS 2023+338 and GRO J0422+32. The flux was continuously growing, from 200 mCrab on Sept. 26 and 550 mCrab on Sept. 27 up to 750 mCrab on Oct. 3 in the band 12-17 keV.”

BINARY MILLISECOND PULSAR IN GRO J0752+17 S. C. Lundgren, A. F. Zepka, and J. M. Cordes, Cornell University,

report: “We announce the discovery of a binary millisecond pulsar in the EGRET source error box for GRO J0752+17. The discovery was made in a pulsar search conducted at Arecibo Observatory targeted at unidentified EGRET high-energy γ-ray source error boxes. The pulsar has a 3.5-ms spin period, an 8.5-hr orbital period, a 430-MHz flux density of 10 mJy, and a dispersion measure of 30 pc cm^{−3}. Based on an estimated success rate of one pulsar per 150 deg^{2} above 3 mJy in past high-latitude surveys, the probability of chance association of a pulsar with one of the observed error boxes is 2 percent. X-ray and γ-ray studies will be necessary to confirm the association of the high-energy source with the radio pulsar.”

1993 October 12 Daniel W. E. Green